OpenStudy (anonymous):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased adult tickets and child tickets.

OpenStudy (anonymous):

5 adult tickets 6 child tickets. 5*9=45 6*4.5=27 27+45=72